### v4

@@ 3.8 Bias Add Unit
There are three Bias_add units for adding biases to the query (), key (), and value () matrices and the outputs of the three feedforward networks. Three units are used because the HLS function definition corresponding to a unit takes arrays of different dimensions as inputs and generates arrays of different dimensions as outputs. One of the units used after feedforward networks contains the ReLU activation function. Algorithm 16, 17 & 13 are the operations of the three units.
3.9 Tiling Technique
As transformer models tend to be large, tiling helps prevent excessive utilization of on-chip memory and computing units. It also ensures that HLS tool can effectively partition arrays, and pipeline or unroll the loops to minimize latency within a short compilation time. Fig. 4(a) describes our unique tiling strategy on MHA. The weight matrices are partitioned into tiles, allowing the BRAMs to be loaded with partial data retrieved from off-chip memory. They are tiled along the second dimension (column of the matrix) only because the first dimension (row of the matrix) is already reduced by the number of heads. Thus, they are loaded () times. The input buffers of each attention head are declared as a two-dimensional array of size (). Therefore, tiling is applied along the column of the matrix and loaded () times. At each iteration, data for only one tile is loaded first. The PEs then perform computations on this data, storing the results in intermediate buffers. These results are also accumulated with those from previous iterations in the next cycle. Consequently, the final output is the cumulative sum of the outputs computed for all tiles.
The FFNs that follow the attention layer are the most time-consuming and resource-consuming layers. The weight matrices of the feedforward network are declared as two-dimensional arrays of size . Thus, they are tiled on both dimensions (row and column of the matrix), and two loops are used to load them iteratively for each tile on each axis. The first FFN module is reused times because both loops iterate times. The second and third FFN modules are reused times. Fig. 4(b) describes our unique tiling strategy on FFN. Results are first accumulated along the columns, followed by accumulation along the rows for all tiles.
3.10 Tile Size Determination
In ADAPTOR, the programmable parameters can be adjusted at runtime, whereas the tile size must be set before synthesis, as it cannot be modified without re-synthesizing the entire hardware. The graph in Fig. 5(a) & (b) illustrates how variations in and impact system frequency (MHz) and latency (normalized to the minimum value), respectively. The number of tiles in MHA () was varied from 6 to 48, for each FFN tile count () which ranged from 2 to 6. The plots indicate that the optimal configuration for achieving the highest frequency and the lowest latency was 24 tiles in MHA and 6 tiles in FFN. This combination would achieve a maximum frequency of 200 MHz.
3.11 Software Control
The parameters such as attention heads, embedding dimension, hidden dimension, sequence length, and the number of encoders and decoders are programmable at runtime in our design. These parameters can be sent to the accelerator from the software using the steps shown in Fig. 6. TNN models are trained using the PyTorch framework, and the appropriate models should be saved as ’.pth’ files. We used pre-trained models available on huggingface [36] on Tesla V100 GPU. These files will then be sent to a Python interpreter to extract the value of the parameters. These values will differ between applications, but our accelerator will not go through the synthesis steps for each one. Only some variables in the software code need to be assigned new values. The Xilinx SDK tool was used to write the software in C++, which runs on the processor. Algorithm 18 briefly describes the software part. Based on the extracted values from the interpreter, the processor generates instructions and control signals for the accelerator, allowing it to activate different parts of the hardware.
3.12 Configuration Registers
The ADAPTOR contains a set of registers accessed by Microblaze CPU using the AXI4-Lite interface. They are used to specify the topology of the TNN during runtime. The registers are described below, along with the corresponding parameters they store.
Sequence: sequence length of inputs.
Heads: number of attention heads.
Layers_enc: number of encoders.
Layers_dec: number of decoders.
Embeddings: dimension of the embedding layer.
Hidden: dimension of the intermediate layers.
Out: number of outputs.
4 Overall System
Fig. 7 shows the complete system design for running ADAPTOR on different FPGA platforms such as VC707 (Virtex-7 xc7vx485tffg1761-2), ZCU102 (Zynq UltraScale+xczu9eg-ffvb1156-2-e MPSoC) and U55C (UltraScale+xcu55c-fsvh2892-2L-e) for our experiments. Both VC707 and ZCU102 boards have onboard DDR3 DRAM memory, while the Alveo U55C contains high-bandwidth memory (HBM). Each design parameter can be programmed during runtime up to a maximum value by the Microblaze CPU. The overall system was designed on Vivado 2022.1.2 design suite. It contains a custom IP block for the TNN accelerator, which was exported from HLS. These Xilinx-AMD tools were run in Intel(R) Xeon(R) Gold 6130 CPU 2.10GHz having 32 cores and 192 GB RAM (host PC). The inputs and weights are taken from the off-chip HBM or DRAM using AXI4 master interfaces [37] when the load instruction from the accelerator controller is received according to demand. The accelerator receives control signals from the processor through an AXI-lite slave interface. The CPU can access the HBMs/DRAMs that are connected to the accelerator. It is used to transfer data to the BRAMs from HBMs. It also sends control signals to the accelerator. The boards are connected to the host PC with USB-JTAG interface and PCIe 3.04 interface. This host can communicate with other IPs except the CPU using the DMA/Bridge Subsystem for PCI Express IP [38] in the system, but PCI communication was not needed in this work. The CPU uses AXI-TIMER[39] to measure the latency, which includes the time between the start and stop signal from the custom IP module. The host connected to JTAG cable[40] displays the results on the terminal using the UARTlite interface[41].
@@ 5 Theoretical Model
The parameters that affect resource utilization and performance of ADAPTOR are the tile size or number of tiles in the attention module, tile size or number of tiles in the feedforward network, number of attention heads, sequence length, embedding dimension, hidden dimension, and number of encoders and decoders when bit width is fixed. The number of DSPs depends on the number of parallel multiplications. They are mostly used in , , , and modules. The number of BRAMs depends on the number of arrays to store data, the BRAM modes applied to them for synthesis, and how they are partitioned using HLS pragmas. An analytical model was developed to establish the relationship of latency and resource utilization with these parameters. This model helps to estimate latency, resource utilization, and the value of the parameters before synthesis.
@@ 5.1 Model for DSP utilization
The design is modular and each module is implemented as a function with loops. Thus, the latency of the modules depends on the loop iteration latency, which in turn depends on the loop pipeline and unrolling pragmas. For the nested loops in the modules, the second loop from the last is pipelined resulting in a complete unroll of the innermost loop. The outermost loop had no pragmas to avoid complicated pipeline depth and resource requirements. Pipelined loop latency (PLL) can be calculated by Eq. 9. If it is enclosed by another loop, then the total latency (TL) will be given by Eq. 10. Here, Trip_count (TC) is the number of iterations of a loop, and the initiation interval (II) is the latency between the initiation of two consecutive iterations. Pipeline_Depth is the latency to finish one iteration. It depends on the sequential and parallel operations executed in one iteration. Different modules of the accelerator can have different Pipeline_Depth (PD). The latency is measured in clock cycles (cc).
@@ 5.2 Latency model for Attention Module
where, PD_L is that includes the time required to establish communication with HBM using AXI master interface (7 cc), read address location (1 cc), load (1 cc), and store (1 cc) data from and to that address, and convert floating point data to fixed point (3 cc) for tasks such as loading all inputs (LI), as well as loading inputs (LIA), biases (LBA) and weights (LWA) for each attention head.
(PD_MHA) equals () plus the time required to load, multiply (2 cc), add (1 cc), and store for computing self-attention (SA) in module. (PD_BA) includes latency associated with loading, adding, and storing operations in bias addition (BA) tasks. (PD_S) equals (), the time required to compute the score (S) in module. (PD_SV) equals in the computation of SV within the module. Equation 19 estimates time for softmax (SM) calculation, which includes exponentiation (4 cc) and division (14 cc). It starts after the module is finished.
@@ 6 Evaluation and Results
ADAPTOR has software programmability to modify the design parameters, including the embedding dimension (), number of heads (h), number of encoder layers (N), sequence length (SL) etc. in runtime from software. They were initially set with fixed values—768, 12, 12, and 64 respectively—based on a variant of BERT[10], which is a transformer model for NLP and the available FPGA resources. The tile sizes cannot be altered after synthesis. Thus, the synthesis was performed with fixed tile sizes of = 64 and = 128. This approach allows ADAPTOR to be synthesized once for a fixed set of resources while retaining the flexibility to support various TNN models as needed.
Fig. 8(a) illustrates the impact of the number of attention heads on system frequency and normalized latency. The latency includes computation time considering that the data loading time is overlapped with computation time. While increasing the number of attention heads is expected to enhance parallelism and reduce latency, the system frequency decreases beyond a certain point, leading to an increase in latency. Optimal performance is observed with 6 to 10 attention heads. Fig. 8(b) shows the increase in resource usage for both DSPs and LUTs as the number of attention heads increases. The high resource utilization is responsible for the reduced frequency.
Fig. 9 illustrates how DSP, LUT, and BRAM utilization is influenced by different combinations of tile sizes (, ). Since the processing modules rely on DSPs for MAC operations, DSPs are the most heavily used and can reach full utilization before BRAMs, making the accelerator computation-bound. As the tile sizes for both the attention and feedforward modules increase, more DSPs are utilized for parallel operations, reducing latency until a drop in frequency occurs.
@@
Fig. 12 presents the roofline model of our accelerator, highlighting its peak performance and memory bandwidth limits. The Memory Bound (blue dashed line) indicates the maximum achievable performance based on the memory bandwidth, which is 103,000 GB/s. Data points to the left of this line are constrained by memory bandwidth. The Compute Bound (red line) represents the peak performance determined by the FPGA’s computational resources, capped at 53 GOP/s. Points below this line indicate underutilization of computational resources. All data points (green, yellow, and purple) fall within the compute and memory-bound regions, meaning none fully utilize the accelerator’s available resources. The yellow square, representing the BERT model with and , achieves the highest performance, being closest to the compute bound. In contrast, the purple star, corresponding to the shallow transformer model with and , exhibits the highest operational intensity but the lowest performance.
@@
Equation 40 below is used to calculate memory bandwidth (BW), where no. of BRAMs 340, BRAM’s width 36 KB, no. of LUTRAMs 129101, LUTRAM’s Width 32 KB. Our previous work [46], mentioned in the cover letter, provides latency and throughput data (53 GOP/s) in Table 1.
[…]
@@
Table 1 compares the performance of our accelerator, ADAPTOR, with other FPGA-based accelerators. Each of these accelerators is optimized for specific TNN models, with some designed specifically for sparse computations. TRAC [48] is the only one that automatically generates accelerator code based on the target FPGA and TNN architecture. Since ADAPTOR was synthesized once with fixed hardware resources and bit width, and implemented on a different platform without sparsity, we evaluated metrics such as throughput (GOPS), power consumption, normalized throughput (GOPS per DSP or GOPS per LUT), and power efficiency (GOPS per watt) for a fair comparison. ADAPTOR achieved 1.9 and 2.25 higher GOPS compared to the accelerators by Qi et al. in [19] and [34], respectively, for a shallow transformer. Its normalized throughput was also higher, indicating more efficient DSP and LUT usage without relying on sparsity, whereas Qi et al. applied block balanced pruning and complex block row storage format. Qi et al.’s four-layer transformer encoder [34] was 1.7 slower and 2 less resource-efficient than ADAPTOR although they applied hierarchical pruning. TRAC [48] used fewer DSPs and LUTs but reported 3.2 higher GOPS and 8.4 higher GOPS/DSP. None of these accelerators applied tiling or described any partitioning scheme to accelerate large models such as BERT, but our work includes those. Tzanos et al.’s work [47] is more comparable to ADAPTOR, as they applied tiling and utilized more resources to achieve speeds 1.6 times faster with GOPS /DSP similar to ADAPTOR.
Fig.13 illustrates the variations in DSP consumption as the tile size of the MHA and FFN layers changes. As DSP utilization increases, GOPS rises up to a certain point. However, as shown in Fig.5, the frequency decreases beyond certain tile sizes. This results in a reduction of GOPS to 30 and 32 for 65% and 70% DSP utilization, respectively.
[…]
@@
[…]
@@
Table 2 compares the experimental data of ADAPTOR with the theoretical results from Sec. 5. Only a few design configurations are presented, with data focused on the computation time for the attention and feedforward modules, as well as the loading time of the attention module for simplicity. Key parameters like sequence length, embedding dimension, and number of heads influence latency. The experimental latency closely matched the theoretical values, with only a 1.8% margin of error. Resource utilization remained consistent across configurations when tile size was fixed. Both analytical and experimental data varied when tile size changed, with deviations of 0.71–4.7% for DSPs and 5.7–74% for BRAMs. The higher deviation from the estimated BRAMs for larger tile sizes occurred because LUTRAMs were used more than BRAMs to maintain high frequency.

### v5

@@ 4.8 Bias Add Unit
Three dedicated Bias_add units are employed to incorporate biases into the query (), key (), and value () matrices, as well as into the outputs of the three feedforward network modules. Separate units are necessary because the corresponding HLS functions operate on arrays of differing dimensions, producing outputs with distinct shapes. One of the units associated with the feedforward networks additionally integrates the ReLU activation function. The operations performed by these units are detailed in Algorithms 16, 17, and 13 of Section 10.
5 System Design and Optimizations
The overall system design for deploying ADAPTOR across multiple FPGA platforms is presented in Fig. 4. Experiments were conducted on three representative devices: the VC707 (Virtex-7 xc7vx485tffg1761-2), ZCU102 (Zynq UltraScale+ xczu9eg-ffvb1156-2-e MPSoC), and the Alveo U55C (UltraScale+ xcu55c-fsvh2892-2L-e). While the VC707 and ZCU102 boards integrate on-board DDR3 DRAM, the Alveo U55C employs high-bandwidth memory (HBM), offering significantly greater throughput for memory-intensive workloads.
Design parameters such as the number of attention heads, embedding dimension, hidden dimension, sequence length, and the number of encoder and decoder layers can be reconfigured at runtime, up to their maximum supported values, via a MicroBlaze softcore processor using the AXI4-Lite interface. These parameters are stored in a set of registers within ADAPTOR to specify the topology of the TNN during runtime from the software. The registers are described in Table 1, along with the corresponding parameters they store. The system architecture was implemented using the Vivado 2022.1.2 design suite. The accelerator itself is encapsulated in a custom IP block generated from high-level synthesis (HLS) and integrated into the larger system. All toolflows from Xilinx-AMD were executed on a host workstation equipped with an Intel(R) Xeon(R) Gold 6130 CPU (2.10 GHz, 32 cores) and 192 GB of RAM. Data movement between the accelerator and external memory is managed through AXI4 master interfaces [41]. Depending on workload demand, the accelerator fetches inputs and weights directly from off-chip HBM or DRAM when instructed by the accelerator controller. Control signals are delivered from the processor to the accelerator via an AXI-Lite slave interface. Additionally, the processor facilitates data transfers between external memory and on-chip BRAMs, while also issuing configuration and synchronization signals to the accelerator.
The boards were connected to the host system through either a USB–JTAG interface or PCIe 3.0 4 link. Although the system includes a DMA/Bridge Subsystem for PCIe IP [42], PCIe-based communication was not utilized in this work. Performance measurements were conducted using the AXI-TIMER [43], which recorded end-to-end latency between the initiation and completion signals of the custom IP. Final results were communicated back to the host via the UARTLite interface [44], with outputs displayed on the terminal connected through the JTAG interface [45].
| Register Name | Description / Stored Parameter |
| Sequence | Sequence length of inputs |
| Heads | Number of attention heads |
| Layers_enc | Number of encoders |
| Layers_dec | Number of decoders |
| Embeddings | Dimension of the embedding layer |
| Hidden | Dimension of the intermediate layers |
| Out | Number of outputs |
The software interface illustrated in Fig. 5 is designed to communicate the programmable parameters mentioned above to the accelerator. To support this, TNN models are trained using the PyTorch framework, with the trained models stored as ’.pth’ files. In our experiments, we utilized publicly available pre-trained models from Hugging Face [46], trained on a Tesla V100 GPU. The software stack processes these files through a Python interpreter, which extracts the relevant parameter values. While these values vary across applications, the accelerator itself does not require re-synthesis for each case. Instead, only a subset of variables within the software must be reassigned to reflect the new configuration. The software, implemented in C++ using the Xilinx SDK and executed on the embedded processor, is summarized in Algorithm 18 of Section 10. Based on the extracted parameters, the processor generates the necessary instructions and control signals to configure the accelerator, thereby enabling selective activation of different hardware components.
The software control overhead varies depending on the execution environment and task complexity. Loading a PyTorch model in Jupyter Notebook typically takes less than 2 seconds for small models and 5 to 20 seconds for large transformer models. Running a Python script, which primarily generates the C code executed in Vitis IDE, generally completes within a few seconds, depending on script complexity and hardware resources. In Vitis IDE, compiling and running the generated C program on the ARM processor (with the bitstream already programmed) requires about 5 to 30 seconds for compilation and 1 to 5 seconds to download and execute the ELF file; rerunning a prebuilt ELF takes less than 2 seconds before execution begins. The FPGA hardware kernels, already synthesized in the bitstream, execute directly with negligible control overhead, and their performance has been reported and compared in the results section.
Transformer models are inherently large, leading to significant demands on both on-chip memory and computational resources. To address these challenges, we adopt a tiling strategy that enables efficient utilization of available hardware resources while maintaining manageable compilation times. Tiling facilitates the effective partitioning of arrays by the HLS tool, which in turn allows loop pipelining and unrolling to reduce computational latency. The proposed tiling approach for multi-head attention (MHA) is illustrated in Fig. 6(a).
In the attention module, the weight matrices are partitioned into tiles, enabling partial data loading from off-chip memory into BRAMs. Tiling is applied along the second dimension (the columns of the matrix), since the first dimension (the rows) is already reduced by the number of attention heads. Consequently, the weight matrices are loaded times. Similarly, the input buffers for each attention head are defined as two-dimensional arrays of size , and tiling is applied along the column dimension. The buffers are replenished times, with one tile being processed at each iteration. During each iteration, the PEs compute results on the loaded tile, store intermediate results in buffers, and accumulate these with the outputs from previous iterations. The final output is thus obtained as the cumulative sum across all tiles. The feedforward networks (FFNs) following the attention layer are the most computationally demanding components of the encoder. Their weight matrices are represented as two-dimensional arrays of size and are tiled along both dimensions (rows and columns). Iterative loading is performed using two nested loops, one for each tiling dimension. As a result, the first FFN module is reused times, since both loops iterate times. The second and third FFN modules are reused times due to their larger dimensions. The tiling strategy for the FFN is summarized in Fig. 6(b), where intermediate results are first accumulated along the columns and subsequently along the rows to produce the final outputs across all tiles.
The tile size must be fixed before synthesis, since changing it would require re-synthesizing the hardware. Figure 7(a) and (b) show how different choices of and affect both system frequency (MHz) and latency (normalized to the minimum value). In these experiments, the number of tiles in MHA () was varied between 6 and 48, while the FFN tile count () ranged from 2 to 6. The results highlight that using 24 tiles for MHA together with 6 tiles for FFN yields the best overall performance, reaching the highest frequency of 200 MHz and the lowest latency.
@@ 6 Theoretical Model
The primary parameters influencing both resource utilization and performance in ADAPTOR include the tile size, or equivalently the number of tiles, in the attention module and the feedforward network, as well as the number of attention heads, sequence length, embedding dimension, hidden dimension, and the number of encoder and decoder layers, assuming a fixed bit width. The utilization of DSPs is largely determined by the degree of parallelism in multiplication operations, with the highest demand observed in the , , , and modules. In contrast, BRAM utilization depends on the number of arrays required for intermediate data storage, the synthesis modes assigned to these memories, and the partitioning strategies specified through HLS pragmas.
To guide design-space exploration, we developed an analytical model that captures the relationship between these architectural parameters and the resulting latency and resource consumption. This model enables designers to predict performance and utilization outcomes, thereby facilitating informed parameter selection prior to full hardware synthesis.
@@ 6.1 Model for DSP utilization
The design follows a modular approach, with each module implemented as a function containing loops. The latency of a module depends on the time taken by its loops, which is affected by loop pipelining and unrolling directives. For nested loops, the second-to-last loop is pipelined, while the innermost loop is fully unrolled. The outermost loop is not modified with pragmas to avoid overly complex pipeline depth and high resource usage. The latency of a pipelined loop (PLL) can be calculated using Eq. 9. If a pipelined loop is inside another loop, the total latency (TL) is given by Eq. 10. Here, the loop trip count (TC) is the number of iterations, and the initiation interval (II) is the time between the start of two consecutive iterations. Pipeline depth is the time needed to complete one iteration, depending on the sequential and parallel operations within it. Different modules can have different pipeline depths (PD). Latency is measured in clock cycles (cc).
@@ 6.2 Latency model for Attention Module
where, PD_L is that includes the time required to establish communication with HBM using AXI master interface (7 cc), read address location (1 cc), load (1 cc), and store (1 cc) data from and to that address, and convert floating point data to fixed point (3 cc) for tasks such as loading all inputs (LI), as well as loading inputs (LIA), biases (LBA) and weights (LWA) for each attention head. (PD_MHA) equals () plus the time required to load, multiply (2 cc), add (1 cc), and store for computing self-attention (SA) in module (Equation 15). (PD_BA) includes latency associated with loading, adding, and storing operations in bias addition (BA) tasks (Equation 16). (PD_S) equals (), the time required to compute the score (S) in module (Equation 17). (PD_SV) equals in the computation of SV within the module (Equation 18). Equation 19 estimates time for softmax (SM) calculation, which includes exponentiation (4 cc) and division (14 cc). It starts after the module is finished.
@@ 7 Evaluation and Results
ADAPTOR supports software-level programmability, allowing modification of key design parameters at runtime. These parameters include the embedding dimension (), number of attention heads (h), number of encoder layers (N), and sequence length (SL). Initially, these parameters were configured with fixed values of 768, 12, 12, and 64, respectively, based on a BERT variant [10], which is a widely used transformer model for natural language processing, and the available FPGA resources. In contrast, the tile sizes are fixed at synthesis and cannot be modified at runtime. Consequently, synthesis was performed with fixed tile sizes of and . This design approach is a key contribution that gives ADAPTOR the ability to retain a single, resource-constrained synthesis configuration while enabling runtime configurability of core transformer parameters so that it can support diverse transformer neural network models without requiring re-synthesis.
Fig. 8(a) presents the effect of varying the number of attention heads on system frequency and normalized latency, where latency accounts for computation time assuming overlap with data loading. While increasing the number of attention heads generally improves parallelism and reduces latency, the system frequency decreases beyond a certain threshold, leading to higher latency. Optimal performance is observed with 6–10 attention heads. Fig. 8(b) illustrates the corresponding increase in DSP and LUT utilization, showing that higher resource usage contributes to reduced system frequency. These results provide a quantitative analysis of the trade-off between parallelism and hardware timing constraints, identify an optimal design point for FPGA-based transformers, and characterize the impact of resource utilization on latency and frequency. Furthermore, the evaluation methodology accounts for overlapped data loading and computation, offering a realistic performance assessment for hardware accelerators.
Fig. 9 illustrates the effect of varying tile sizes (, ) on the utilization of DSP, LUT, and BRAM. Since processing modules rely on DSPs for multiplication–accumulation (MAC) operations, DSPs represent the most widely used resource and can reach saturation before BRAMs, rendering accelerator computation bound. Increasing the tile sizes for both the attention and feedforward modules results in higher DSP utilization, which enables greater parallelism and reduces latency until the system frequency begins to decline. This analysis is the characterization of resource–performance trade-offs, showing that tile size selection directly determines when the accelerator becomes computation-bound or frequency-limited, thus guiding optimal design choices for FPGA-based transformer implementations.
@@
Fig. 12 presents the roofline model of ADAPTOR, highlighting its peak performance and memory bandwidth limits. The Memory Bound (blue dashed line) indicates the maximum achievable performance based on the memory bandwidth, which is 103,000 GB/s. Data points to the left of this line are constrained by memory bandwidth. The Compute Bound (red line) represents the peak performance determined by the FPGA’s computational resources, capped at 53 GOP/s. Points below this line indicate underutilization of computational resources. All data points (green, yellow, and purple) fall within the compute and memory-bound regions, meaning none fully utilize the accelerator’s available resources. The yellow square, representing the BERT model with and , achieves the highest performance, being closest to the compute bound. In contrast, the purple star, corresponding to the shallow transformer model with and , exhibits the highest operational intensity but the lowest performance.
@@
Equation 40 below is used to calculate memory bandwidth (BW), where no. of BRAMs 340, BRAM’s width 36 KB, no. of LUTRAMs 129101, LUTRAM’s Width 32 KB. Our previous work [51] calculated latency and throughput data (53 GOP/s).
| Accelerator | DSP | LUT | GOPS | Power | (GOPS/DSP) | (GOPS/LUT) | GOPS/ | Method | Sparsity |
| (W) | 1000 | 1000 | Power | ||||||
| Network #1 | Shallow Transformer | ||||||||
| Qi et al. [19] | 3572 (52%) | 485 k (41%) | 14 | – | 3.92 | 0.03 | – | HLS | 80% |
| Qi et al. [34] | 5040 (74%) | 908 k (76%) | 12 | – | 2.38 | 0.013 | – | 86% | |
| ADAPTOR | 3612 (40%) | 391 k (30%) | 27 | 11.8 | 7.47 | 0.069 | 2.28 | 0% | |
| Network #2 | Custom Transformer Encoder | ||||||||
| Qi et al. [34] | 4145 (60%) | 937 k (79%) | 75.94 | – | 18 | 0.08 | – | HLS | 0% |
| ADAPTOR | 3612 (40%) | 391 k (30%) | 132 | 11.8 | 37 | 0.34 | 11 | ||
| Network #3 | BERT | ||||||||
| Tzanos et al. [52] | 5861 (85%) | 910 k (77%) | 65.7 | – | 11.2 | 0.07 | – | 0% | |
| TRAC [39] | 1379 (80%) | 126 k (55%) | 128 | – | 93 | 1.01 | – | – | |
| ADAPTOR | 3612 (40%) | 391 k (30%) | 40 | 11.8 | 11 | 0.10 | 3.39 | 0% | |
@@
Table 2 compares the performance of our accelerator, ADAPTOR, with other FPGA-based accelerators. Each of these accelerators is optimized for specific TNN models, with some designed for sparse computations. TRAC [39] is the only one that automatically generates accelerator code based on the target FPGA and TNN architecture. Since ADAPTOR was synthesized once with fixed hardware resources and bit width, and implemented on a dense model without sparsity, we evaluated throughput (GOPS), power consumption, normalized throughput (GOPS per DSP or GOPS per LUT), and power efficiency (GOPS per watt) for a fair comparison. ADAPTOR achieved 1.9 and 2.25 higher GOPS compared to the accelerators by Qi et al. in [19] and [34], respectively, for a shallow transformer. Its normalized throughput was also higher, indicating more efficient DSP and LUT usage without relying on pruning, whereas Qi et al. employed block balanced pruning and block row storage. Qi et al.’s four-layer transformer encoder [34] was 1.7 slower and 2 less resource-efficient than ADAPTOR even with hierarchical pruning. TRAC [39] consumed fewer DSPs and LUTs but reported 3.2 higher GOPS and 8.4 higher GOPS/DSP. None of these accelerators incorporated tiling or partitioning schemes to support large models such as BERT, which our design explicitly addresses. Tzanos et al. [52] applied tiling and used more resources, achieving 1.6 higher speed with GOPS/DSP comparable to ADAPTOR.
Although ADAPTOR utilizes over 3,000 DSPs at 200 MHz, the measured throughput is significantly lower than the theoretical peak of 1200 GOPS. This underutilization arises from several factors: (i) sequential execution of certain submodules (e.g., , , , , , , softmax, layer normalization etc.) that prevents simultaneous activation of all functional modules, and (ii) control dependencies that limit pipelining across nested loops. As a result, a portion of the available DSPs remain idle at different stages of execution, reducing overall efficiency. Fig. 13 illustrates how GOPS scales with DSP consumption as the tile sizes of the MHA and FFN layers increase. While larger tiles increase DSP utilization and improve throughput, the system frequency drops beyond certain tile sizes (Fig. 7), leading to diminishing returns and even a reduction of GOPS to 30 and 32 for 65% and 70% DSP utilization, respectively. This analysis highlights the fundamental trade-off between resource utilization, frequency, and achievable throughput in FPGA-based transformer accelerators.
| Method | Sequence | Embedding | Number | Tile | Tile | DSPs | BRAMs | Frequency | Latency (ms) | ||
| Length | Dimension | of | Size | Size | 18k | (MHz) | Attention | Load | FFN | ||
@@
| (SA) | (LWA) | (FFN1) | |||||||||
| Analytical | 64 | 768 | 8 | 64 | 128 | 3784 | 2375 | 200 | 0.052 | 0.037 | 0.082 |
| Experimental | 3612 | 2246 | 0.053 | 0.038 | 0.084 | ||||||
| Analytical | 128 | 768 | 8 | 64 | 128 | 3784 | 2375 | 0.103 | 0.037 | 0.165 | |
| Experimental | 3612 | 2246 | 0.106 | 0.038 | 0.168 | ||||||
| Analytical | 64 | 512 | 8 | 64 | 128 | 3784 | 2375 | 0.042 | 0.025 | 0.055 | |
| Experimental | 3612 | 2246 | 0.043 | 0.026 | 0.056 | ||||||
@@
Table 3 presents a comparison between the experimental results of ADAPTOR and the theoretical predictions derived in Section 6. For clarity, only a subset of design configurations is reported, focusing on the computation time of the attention and feedforward modules as well as the loading time of the attention module. Latency is primarily influenced by parameters such as sequence length, embedding dimension, and number of attention heads. The measured latency closely aligned with the theoretical estimates, with an average deviation of only 1.8%. Resource utilization remained stable across configurations with fixed tile sizes, whereas variations in tile size led to corresponding changes in both analytical and experimental values. The deviations were relatively small for DSPs (0.71–4.7%), but larger for BRAMs (5.7–74%), particularly at larger tile sizes. The latter discrepancy arises because LUTRAMs were increasingly used in place of BRAMs to sustain higher operating frequency, thereby reducing the accuracy of BRAM utilization estimates.